Solve the problem by all three methods. Then, write a paragraph comparing and contrasting each method.

            

Prepare 0.1 g of an IM medication from a strength of 200 mg per mL.
Answer #1

200mg of the medicine is in 1ml of the medication.

200mg =0.2g

dose in hand = 0.2 g

dose prescribed = 0.1 g

amount in hand = 1ml

How much ml of the medicine contain 0.1 g of the medicine?

dose prescribed x amount in hand / dose in hand = 0.1 x1 /0.2 = 1/2 = 0.5 ml